If 2/3 of a number is 10, then what is 1.75 times of that number?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

From the question it is given that,
2/3 of a number is 10.

Let us assume the number be ‘P’.Then,

2/3 of P = 10

2/3 × P = 10
By cross multiplication we get,

P = 10 × 3/2 P= 5 × 3
P = 15

So, the number is 15
Again it is given in question that, 1.75 times of that number = ?

= 1.75 of 15
= 1.75 × 15
= 26.25


Thanks